OpenStudy (anonymous):

how do you divide 7743 by 4 without using long division? please show me every step

OpenStudy (lgbasallote):

you can break it up... \[\implies \frac{7740 + 3}{4} \implies \frac{7740}{4} + \frac 34 \implies 1935 + 0.75\] is that what you mean?

OpenStudy (anonymous):

oh thankss sooo muchh

OpenStudy (lgbasallote):

welcome